Write each conic sectioning standard form. Identify the conic section.
3. x^2+y^2-2x-24=0
4. X^2+4y^2+24y+32=0
5. 3x^2-5x+3y^2-8=0

Answer by lwsshak3(11628)

Write each conic sectioning standard form. Identify the conic section.
3. x^2+y^2-2x-24=0
4. X^2+4y^2+24y+32=0
5. 3x^2-5x+3y^2-8=0
..
3. x^2+y^2-2x-24=0
(x^2-2x+1)+(y^2-0)=24+1=25
(x-1)^2+y^2=25
This equation is a circle with center at (1,0) with radius=5
..
4. X^2+4y^2+24y+32=0
x^2+4(y^2+6y+9)=-32+36=4
x^2+4(y+3)^2=4
x^2/4+(y+3)^2=1
This equation is an ellipse with center at (0,-3) with a horizontal major axis
..
5. 3x^2-5x+3y^2-8=0
3(x^2-5x/3+25/36)+3y^2=8+25/12=96/12+25/12=121/12
3(x-5/6)^2+3y^2=121/12
divide by 3
(x-5/6)^2+y^2=121/36
This equation is a circle with center at (5/6,0) with radius=sqrt(121/36)